OpenStudy (anonymous):

factor this expression. Check by multiplying factors. 3ab+a+3b^2+b

OpenStudy (anonymous):

= 3b(a - b) - (a - b) = (a-b)(3b - 1)

OpenStudy (koikkara):

hmm... @bpscrewer Sure ?

OpenStudy (anonymous):

See, I was thinking 3b(a+b)

OpenStudy (anonymous):

i believe so?

OpenStudy (koikkara):

Good, its how you solve... @Instagram !!

OpenStudy (anonymous):

so am i wrong?

OpenStudy (anonymous):

Okay, so first term done. Now for the second term.. I'm not sure how to use the exponent or variable.

OpenStudy (anonymous):

\(\Large 3ab+a+3b^2+b=(3ab+3b^2)+(a+b)\\\Large =3b(a+b)+(a+b)=(a+b)(3b+1)\)

OpenStudy (koikkara):

@bpscrewer = 3b(a - b) - (a - b) = 3ba-3b^2-a=b = .... it goes on... much complicated.. here we get -a !

OpenStudy (koikkara):

@@bpscrewer Hope your doubt is clear.... as in question it has +a not -a !

OpenStudy (anonymous):

I'm so lost..

OpenStudy (anonymous):

yea i kinda am too @Instagram

OpenStudy (koikkara):

@Instagram Hmm, i was explaining to @bpscrewer were he/she went wrong !! \(\color{darkgreen}{~''Nice~To~Meet~You~!!"}\)

OpenStudy (anonymous):

Alrighty.

OpenStudy (anonymous):

So is the answer (a+b)(3b+1)

OpenStudy (koikkara):

ref: @JungHyunRan answer @Instagram ! She already did it for you...!

OpenStudy (anonymous):

Oh, I didn't see that. Thanks!

OpenStudy (koikkara):

\(\color{red}{~''Bingo~!!~There~You~Go~!!"}\) Have A nice Day Guys !!